1.      (i) Use normal approximation to a Binomial Distribution (X) having n=500 trials and p = 0.025 for finding out:

(a)   P(X>10);                              (b) P(X<18)       (c) P(X>21)               (d) P(9<X<14)   

        

    


1
Expert's answer
2021-07-18T17:39:14-0400

"p=0.025 \\

\mu= np \\

mu= 50 \times 0.025= 12.5 \\

\sigma= \sqrt{np(1-p)} \\

\sigma= \sqrt{12.5(1-0.025)} = 3.49"

With continuity correction factor

(a)

"P(X>10) = P(X>10.5) \\

= 1 -P(X<10.5) \\

= 1 -P(Z< \frac{10.5-12.5}{3.49}) \\

= 1 -P(Z< -0.5730) \\

= 1 -0.2833 \\

=0.7167"

(b)

"P(X<18) = P(X<17.5) \\

= P(Z < \frac{17.5-12.5}{3.49}) \\

= P(Z< 1.4326) \\

= 0.9240"

(c)

"P(X>21) = P(X>21.5) \\

= 1 -P(X<21.5) \\

= 1 -P(Z< \frac{21.5-12.5}{3.49}) \\

= 1 -P(Z< 2.5787) \\

= 1 -0.9950 \\

= 0.005"

(d)

"P(9<X<14) = P(8.5<X<14.5) \\

= P(X<14.5) -P(X<8.5) \\

=P(Z< \frac{14.5-12.5}{3.49}) -P(Z< \frac{8.5-12.5}{3.49}) \\

= P(Z< 0.5730) -P(Z< -1.1461) \\

= 0.7166 -0.1258 \\

= 0.5908"